A test car accelerates from 0 to 60mph in 4.9 seconds. what is the average acceleration in stadard SI units. Assume1609 meters in one mile.
Answer #1

\large 1 mile = 1609m

\large 60mph=26.8224m/s

\large u=0m/s,v=26.8m/s,t=4.9s,t=?

\large a=\frac{v-u}{t}=\frac{(26.8m/s-0m/s)}{4.9s}

\large a=5.469ms^{-2}

average acceleration \large a=5.47ms^{-2}